OpenStudy (itsbribro):

Which statement describes the graph and system of equations? 3y-x=12 x-3y=6 A. coincident B. consistent and independent C. inconsistent D. not enough info

OpenStudy (itsbribro):

@kah.x @Destinyyyy @whpalmer4 @Agl202

OpenStudy (itsbribro):

i cant post the graph. it won't let me

OpenStudy (itsbribro):

yes

OpenStudy (agl202):

I would say it is B. consistent and independent. Points that work in each equation...consistent. When a system is "dependent," it means that all points that work in. One of them also work in the other one. Graphically, this means that. One line is lying entirely on top of the other one, so that if you graphed both, you would really see only one line on the graph. Since they are imposed on top of each other. One of them totally depends on the other one. So, B. is what I say.

OpenStudy (itsbribro):

ohhhhh okay now i kinda understand lol

OpenStudy (itsbribro):

i think its inconsistent :/

OpenStudy (itsbribro):

but thank you for your help :) i appreciate it

OpenStudy (whpalmer4):

This is an inconsistent system of equations. To quote the wikipedia article on systems of linear equations: A linear system is inconsistent if it has no solution, and otherwise it is said to be consistent. When the system is inconsistent, it is possible to derive a contradiction from the equations, that may always be rewritten such as the statement 0 = 1. \[3y-x=12\]\[x-3y=6\] Look what happens if we solve the second equation for \(x\) and substitute that into the first: \[x-3y=6\]\[x=3y+6\]\[3y-x=12\]\[3y-(3y+6)=12\]\[3y-3y-6=12\]\[-6=12\]